# Condensate Size Control by Charge Asymmetry

### v1

@@ Abstract
Biomolecular condensates are complex droplets comprising many different types of molecules that interact using various mechanisms. Condensation is often driven by short-ranged attraction, but net charges can also mediate long-ranged repulsion. Using molecular dynamics simulations and an equilibrium field theory, we show that such opposing interactions can suppress coarsening so that many droplets of equal size coexist at equilibrium. This size control depends strongly on the charge asymmetry between constituents, while the strength of the short-ranged attractions has a weak influence. Our work reveals how electrostatic effects control droplet size, which is relevant for understanding biomolecular condensates and creating synthetic patterns in chemical engineering.
Biomolecular condensates are complex droplets that are key for various cellular functions [1]. They typically comprise many different biomolecules, including nucleic acids and proteins, which can be charged [2]. The resulting electrostatic interactions are crucial for direct interactions, which can drive phase separation and condensate formation [3, 4]. Moreover, droplets can form by complex coacervation of oppositely charged polymers [5], which is crucial in various biological systems [6, 7]. Simulations [8], experiments [9], and theory [10, 10] suggest that complex coacervation is driven primarily by entropic effects. Yet, long-ranged electrostatic interactions can also be significant in charged droplets, which have been observed experimentally [6] and are predicted for systems with charge asymmetries [11]. Typical condensates might thus exhibit a combination of short-ranged attraction driving phase separation and long-ranged repulsion from net charges.
@@
Previous theoretical work suggests that a net charge of droplets affects their interfaces [12, 13], reduces coarsening [14, 15, 16], and can potentially lead to fission [17, 18]. Even though simulations demonstrate reduced surface tensions [19], coalescence is suppressed by an energy barrier [15]. Simulations also find that heterogeneous charges can lead to multiphasic coacervates [20]. While phase separation tendency of polyelectrolytes increases with overall charge [21], a large charge asymmetry can suppress phase separation completely [19]. These observations indicate that net charge fundamentally affects phase separation of oppositely charged polymers, which might be particularly relevant in condensates, but the precise mechanism is elusive.
@@
Figure 1: MD simulations show smaller droplets for larger charge asymmetry. (a) Schematic of simulations set-up, where a system of two oppositely charged polyions is neutralized with counterions, replicated, and simulated in the $NVT$ ensemble (i.e., constant number of particles, volume, and temperature). (b) Snapshots from three simulations at various charge pairings of the polycation ($z_{\mathrm{C}}$, red) and polyanion ($z_{\mathrm{A}}$, blue) showing that larger charge asymmetry leads to smaller droplets. For clarity, counterions are shown 99% transparent.
@@
To investigate the influence of charge asymmetry on phase separation, we first performed molecular dynamics (MD) simulations of a system consisting of polycations $C$, polyanions $A$, and the respective counterions, $e^{-}$ and $e^{+}$ (Fig. 1a). For simplicity, we kept the length of the polyions fixed at ten times the size of an ion and instead varied the respective charge numbers per monomer, $z_{\mathrm{C}}$ and $z_{\mathrm{A}}$, to study different charge asymmetries $\Delta z=z_{\mathrm{C}}+z_{\mathrm{A}}$. We ensured charge neutrality of the entire system by compensating each charge on polyions by respective counterions. We modeled the direct interactions between oppositely charged polyions by attractive Lennard-Jones interactions and used excluded volume interactions between all other species. Additional details are described in the Appendix. The snapshots of equilibrated MD simulations shown in Fig. 1b demonstrate that polyions form dense droplets surrounded by a dilute phase. Importantly, we find smaller droplets for larger charge asymmetry, indicating that droplet size is controlled and multiple droplets can stably coexist.
@@
To reveal the physical mechanism of the observed droplet size regulation, we next describe the system using a continuous field theory. Here, the state of the incompressible, isothermal mixture is captured by the volume fraction fields $\phi_{\mathrm{C}}(\boldsymbol{r})$, $\phi_{\mathrm{A}}(\boldsymbol{r})$, $\phi_{\mathrm{e^{+}}}(\boldsymbol{r})$, and $\phi_{\mathrm{e^{-}}}(\boldsymbol{r})$ of the charged species. The remaining fraction, $\phi_{\mathrm{S}}=1-\sum_{i}\phi_{i}$ (here and below sums are over the four species $C$, $A$, $e^{+}$, and $e^{-}$), is filled by an inert solvent $S$. The system’s equilibrium is governed by the minimum of the total free energy $F=F_{\mathrm{FH}}+F_{\mathrm{int}}+F_{\mathrm{el}}$, where the terms respectively capture local, interfacial, and long-ranged electrostatic interactions akin to Ref. [13]. We approximate local interactions using the Flory–Huggins model [22, 23],
$F_{\mathrm{FH}}=\frac{k_{\mathrm{B}}T}{\nu}\!\int\bigg{[}\chi\phi_{\mathrm{A}}\phi_{\mathrm{C}}+\phi_{\mathrm{S}}\ln\phi_{\mathrm{S}}+\sum_{i}\frac{\phi_{i}}{l_{i}}\ln\phi_{i}\biggl{]}\mathrm{d}V\;,$ (1)
@@
where integrals are over the system of volume $V$, $k_{\mathrm{B}}T$ is the relevant thermal energy scale, and $\nu$ denotes the molecular volume, where solvent molecules, ions, and monomers of the polyions have the same volume for simplicity. In the integrand, the first term proportional to the Flory parameter $\chi<0$ accounts for attraction between polycations $C$ and polyanions $A$, whereas the other terms capture translational entropies. To mimic the MD simulations, we chose the respective chain lengths $l_{i}$ as $l_{e^{+}}=l_{e^{-}}=1$ and $l_{\mathrm{C}}=l_{\mathrm{A}}=10$. In contrast, the interfacial energy is described by
@@
where $\varepsilon$ is the dielectric constant, which we assumed to be spatially invariant. The charge numbers $z_{i}$ represent the degree of ionization of each monomer, in units of the elementary charge $e$, so the charge asymmetry is $\Delta z=z_{\mathrm{C}}+z_{\mathrm{A}}$. Taken together, the equilibrium state thus depends on the local attraction ($\chi$), the interfacial penalty ($\kappa$), and the charge numbers ($z_{\mathrm{C}}$ and $z_{\mathrm{A}}$).
@@
Figure 2: Field theory predicts droplet-like periodic patterns. Periodic profiles of charge density $\rho_{i}=z_{i}\phi_{i}$ for all species $i$ and total charge density $\rho=\sum_{i}\rho_{i}$ as a function of distance for weak (left) and strong (right) charge asymmetry $\Delta z=z_{\mathrm{C}}+z_{\mathrm{A}}$. Parameters are $z_{\mathrm{A}}=-0.2$, $l_{\mathrm{C}}=l_{\mathrm{A}}=10$, $l_{e^{+}}=l_{e^{-}}=1$, $\bar{\phi}_{\mathrm{C}}=\bar{\phi}_{\mathrm{A}}=0.1$, $\kappa=10\,\ell_{\mathrm{B}}^{2}$, and $\chi=-5$, with Bjerrum length $\ell_{\mathrm{B}}=e^{2}/(\varepsilon k_{\mathrm{B}}T)$.
@@
To mimic the MD simulations, we impose constant average fractions $\bar{\phi}_{i}=V^{-1}\int\phi_{i}\mathrm{d}V$ with $\bar{\phi}_{\mathrm{C}}=\bar{\phi}_{\mathrm{A}}=0.1$. To ensure global charge neutrality, we set $z_{e^{+}}=-z_{e^{-}}=1$ and adjust the fraction of counterions, $\bar{\phi}_{e^{-}}=z_{\mathrm{C}}\bar{\phi}_{\mathrm{C}}$ and $\bar{\phi}_{e^{+}}=-z_{\mathrm{A}}\bar{\phi}_{\mathrm{A}}$. Based on the results shown in Fig. 1b, we expect that equilibrium states exhibit droplets of a well-defined size, which corresponds to periodic profiles in the field theory. To describe such states, we for simplicity consider a one-dimensional, periodic system of length $L$, and minimize $F$ by varying $L$, $\psi(x)$, and $\phi_{i}(x)$. Here, we allow for coexisting phases with different periods and we impose mass conservation, incompressibility, and charge neutrality within each phase using Lagrange multipliers, akin to Ref. [25] and described in the Appendix. The resulting equilibrium profiles shown in Fig. 2 indicate that the field theory indeed predicts periodic patterns, where regions of large polyion density correspond to the droplets and the surrounding region enriched in counterions represent the dilute phase. However, the counterions do not neutralize the system everywhere (black line), which we confirmed in MD simulations (Fig. S1 in the Appendix). These net charges are reminiscent of the double layer structure reported for complex coacervates [26] and indicate that electrostatics are crucial for droplet size control [11]. Indeed, we observe larger periods $L$ for smaller charge asymmetry (Fig. 2), consistent with the larger droplets in the MD simulations (Fig. 1b). There are also subtle differences between the two panels: For small charge asymmetry (Fig. 2a), the charge density $\rho$ exhibits a dip inside the droplet and reaches neutrality ($\rho=0$) in the dilute phase. Both effects are absent for larger charge asymmetry (Fig. 2b), suggesting that these two cases are qualitatively different.
@@
Figure 3: Field theory predicts transitions to patterned phases. (a) Phase diagram as a function of the charge numbers of polycations ($z_{\mathrm{C}}$) and polyanions ($z_{\mathrm{A}}$) revealing parameter regions with the coexistence of two homogeneous phases (H+H) and patterned phases (P) for strong attraction ($\chi=-5$). The colored markers correspond to the two panels in Fig. 2. (b) Phase diagram for weak attraction ($\chi=-4.5$) including regions with a single homogeneous phase (H). The remaining columns show respective periods $L$ and amplitudes $A_{\mathrm{C}}=\max(\phi_{\mathrm{C}})-\min(\phi_{\mathrm{C}})$ of the polycation profile. (c) Schematic explanation of the origin of the patterned phase. The initial phase-separated charge-neutral state in (i) is destabilized due to net charges if charge asymmetry is enabled in (ii). An equilibrium state can either be reached by restoring charge neutrality in (iii) or by breaking the phases into smaller droplets in (iv). (d) Pattern length scale $L$ as a function of interaction parameter $\chi$ (left, interfacial parameter $\kappa=10$) and $\kappa$ (right, various $\chi$) for various $z_{\mathrm{C}}$ at $z_{\mathrm{A}}=-0.2$. As $|\chi|$ decreases, the patterned phase transitions to macrophase separation (stars) or a homogeneous state (open disk). Dashed lines indicate the most unstable length scale predicted by the linear stability analysis in the Appendix. (a–d) Additional parameters are given in Fig. 2.
@@
We next analyze in detail how the periodic patterns depend on the charge numbers $z_{\mathrm{C}}$ and $z_{\mathrm{A}}$. Fig. 3a shows that periodic patterns only appear in a restricted parameter region (yellow region P). In particular, the patterned phase does not exist for symmetric mixtures (diagonal line, $\Delta z=0$). Instead, the system exhibits macrophase separation (blue H+H region) if $\chi$ is sufficiently negative, and this region becomes larger for weaker interactions (Fig. 3b and Fig. S6 in the Appendix.) Moreover, the system stays homogeneous for very asymmetric mixtures (teal H region). Taken together, these observations suggest periodic patterns emerge for sufficiently strong attraction (large $-\chi$) and intermediate charge asymmetry $|\Delta z|$ between polyions.
@@
We can understand the emergence of patterned phases starting from the charge-balanced state ($\Delta z=0$, Fig. 3c, subpanel (i). In this case, the polyions form a dense macrophase while the small molecules ($e^{+}$, $e^{-}$, and $S$) accumulate in the corresponding dilute phase. Here, we have $F=F_{\mathrm{FH}}$ since $F_{\mathrm{el}}=0$ and the energy $F_{\mathrm{int}}$ of the single interface is negligible in a thermodynamically large system. Switching on charge asymmetry ($\Delta z\neq 0$, subpanel (ii) destabilizes this state because the macrophases acquire a net charge, so the electrostatic energy $F_{\mathrm{el}}$ diverges. To minimize the total free energy $F$, the state can change in two fundamentally different ways: it can either restore charge neutrality (subpanel iii) or form a patterned phase (subpanel iv). In the first alternative, counterions move into the dense phase (maintaining macrophase separation, subpanel (iii), which restores charge neutrality and maintains $F_{\mathrm{el}}=F_{\mathrm{int}}=0$. However, the small counterions decrease the concentration of polyions, thus reducing their contact and increasing $F_{\mathrm{FH}}$. In the second alternative, the dense phase is broken up to form a pattern of finite length scale $L$ (subpanel iv). In the simplest case, the polyions maintain a high concentration, implying that $F_{\mathrm{FH}}$ changes negligibly. However, the electrostatic energy $F_{\mathrm{el}}$ is now finite, and the breakup creates many interfaces, so both $F_{\mathrm{el}}$ and $F_{\mathrm{int}}$ are positive. Taken together, both alternatives increase $F$ compared to the neutral initial case, but breakup into the patterned phase is favored for strong attraction (large negative $\chi$).
@@
The preceding argument also makes qualitative predictions for the pattern period $L$ as a proxy for droplet size. Assuming $F_{\mathrm{FH}}$ stays roughly constant in the patterned phase, $L$ will depend only weakly on $\chi$ and is instead governed by the minimum of $F_{\mathrm{int}}+F_{\mathrm{el}}$. Generally, $F_{\mathrm{int}}$ decreases with increasing $L$ (since there are fewer interfaces per unit length), whereas $F_{\mathrm{el}}$ increases with $L$ (since opposite charges are separated further for larger $L$, i.e., exchanging the two central slabs in Fig. 3c(iv) costs electrostatic energy). These arguments predict that $L$ increases with larger $\kappa$ (increasing $F_{\mathrm{int}}$), whereas $L$ decreases with larger $|\Delta z|$ (increasing $F_{\mathrm{el}}$). We present a more quantitative argument in the Appendix, which shows that at the free energy minimum the two energies have opposing dependencies on $L$, $\mathrm{d}F_{\mathrm{int}}/\mathrm{d}L=-\mathrm{d}F_{\mathrm{el}}/\mathrm{d}L$, leading to a trade-off where $F_{\mathrm{int}}=F_{\mathrm{el}}$ at the minimum. We thus generally expect that $L$ increases with larger $\kappa$ and smaller $|\Delta z|$.
@@
We test our predictions by analyzing the patterned phases obtained from the field theory. Fig. 3b shows that $L$ indeed decreases for larger $|\Delta z|$ (away from the symmetry line), although the patterns completely vanish for large $|\Delta z|$ where the homogeneous state exhibits the lowest free energy. The right-most panel in Fig. 3b shows that the amplitude of $\phi_{\mathrm{C}}(x)$ decreases with increasing $|\Delta z|$, implying that polyions become less concentrated. This is not captured by the simple argument above, suggesting that the system compromises the segregation of polyions (implying larger $F_{\mathrm{FH}}$) for a lower electrostatic energy $F_{\mathrm{el}}$. Indeed we find that $F_{\mathrm{el}}$ exhibits a non-monotonous dependency on $|\Delta z|$ (Fig. S11 in the Appendix), and the patterns cease when $F_{\mathrm{el}}$ vanishes. Concomitantly, the amplitude approaches zero (Fig. 3b), suggesting that the transition to the homogeneous phase is continuous.
@@
To also test the influence of the interaction strength $\chi$ and the interfacial parameter $\kappa$, we performed additional minimizations. Fig. 3d confirms a weak influence of $\chi$, whereas $L$ increases significantly with $\kappa$. Note that the length scale predicted from the most unstable perturbation around the homogenous state (dashed line in Fig. 3d; details in the Appendix) generally does not predict the equilibrium periods. However, both values coincide at the transition of the patterned phase to the homogeneous state (open circles) and the scaling is consistent with mean-field theory (Fig. S5 in the Appendix), supporting that this is a continuous phase transition. Taken together, the numerical solutions of the field theory confirm that $L$ (and thus the droplet size) increases with larger $\kappa$ and smaller $|\Delta z|$, while $\chi$ has a minor influence beyond driving phase separation in the first place.
@@
Figure 4: MD simulations corroborate field theory. (a) Simulations at various charge numbers ($z_{\mathrm{C}}$, $z_{\mathrm{A}}$) demonstrate that high charge asymmetries result in smaller average droplet sizes (normalized by total number of polyion chains in the system). (b) Normalized droplet size as a function of attraction strength $\epsilon_{\mathrm{LJ}}$. Decreasing $\epsilon_{\mathrm{LJ}}$ mildly increases or decreases droplet size depending on charge asymmetry.

### v2

# Condensate Size Control by Net Charge

@@ Abstract
Biomolecular condensates are complex droplets comprising diverse molecules that interact using various mechanisms. Condensation is often driven by short-ranged attraction, but net charges can also mediate long-ranged repulsion. Using molecular dynamics simulations and an equilibrium field theory, we show that such opposing interactions can suppress coarsening so that many droplets of equal size coexist at equilibrium. This size control depends strongly on the charge asymmetry between constituents, while the strength of the short-ranged attractions has a weak influence. Essentially, droplets expel ions, so they cannot screen electrostatics effectively, implying droplets acquire a net charge and cannot grow indefinitely. Our work reveals how electrostatic effects control droplet size, which is relevant for understanding biomolecular condensates and creating synthetic patterns in chemical engineering.
Biomolecular condensates are complex droplets that are key for numerous cellular functions [1]. They typically comprise diverse biomolecules, including nucleic acids and proteins, and form by phase separation due to various short-ranged interactions. On microscopic length scales, typical interactions include $\uppi$$\uppi$-stacking, cation–$\uppi$ interactions, hydrogen bonding, and hydrophobic interactions [2, 3]. Moreover, many of the involved molecules are charged to different degrees [3], implying electrostatic interactions. These electrostatic interactions can contribute to the short-ranged interactions leading to phase separation. An example are complex coacervates, where oppositely-charged polyions attract each other and phase separate from a solvent [4, 5, 6, 7, 8]. Electrostatics also induce interactions between macroscopic regions with net charge, although such long-ranged interactions are typically screened by mobile salt ions that neutralize charges on macromolecules in dilute electrolytes like the cytosol [9]. However, screening might be suppressed if phase separation affects the distribution of ions. Taken together, short-ranged interactions are crucial for forming biomolecular condensates, but the influence of long-ranged electrostatic effects is unclear.
@@
Previous theoretical work suggests that long-ranged electrostatic effects are relevant in phase separating systems. For instance, a net charge of droplets affects their interfaces [10, 11], reduces coarsening [12, 13, 14], and can potentially lead to fission [15, 16]. In particular, reduced surface tension [17] and an energy barrier suppressing coalescence [13] slow down coarsening. Simulations also find that heterogeneous charges can lead to multiphasic coacervates [18], and while polyelectrolytes with more charges phase separate more strongly [19], a large charge asymmetry can suppress phase separation completely [17]. These observations indicate that net charge fundamentally affects phase separation of macromolecules, but the precise mechanism, and the role played by ions, is elusive.
@@
Figure 1: MD simulations show smaller droplets for larger charge asymmetry. (a) Schematic of simulations set-up, where a system of two oppositely charged polymers is neutralized with an excess of ions, replicated, and simulated in the $NVT$ ensemble (constant number of particles, volume, and temperature). (b) Snapshots from three simulations at various charges of the positively-charged ($z_{P^{+}}$, red) and negatively-charged polymers ($z_{P^{-}}$, blue) showing that larger charge asymmetry leads to smaller droplets. For clarity, ions are shown 99% transparent.
@@
To investigate phase separation of charged macromolecules in the presence of ions, we first performed molecular dynamics (MD) simulations of a system consisting of positively-charged polymers $P^{+}$, negatively-charged polymers $P^{-}$, and related ions, $e^{+}$ and $e^{-}$ in the NVT ensemble with an implicit solvent modeled using a Langevin thermostat (Fig. 1a). We induced phase separation using an attractive Lennard-Jones interaction between $P^{+}$ and $P^{-}$, while all other species exhibit excluded volume interactions. For simplicity, we kept the length of the polymers fixed at ten times the size of an ion, and initialize the system such that $P^{+}$ and $P^{-}$ each occupy $10\%$ of the volume. Electrostatic interactions are given by Coulomb potentials between all species, which have charge numbers $z_{i}$. In particular, we varied the respective charge numbers $z_{P^{+}}>0$ and $z_{P^{-}}<0$ per monomer, whereas the ions have $z_{e^{+}}=1$ and $z_{e^{-}}=-1$. To have comparable electrostatic effects, we added as many positive ions $e^{+}$ as there are charges on the negatively-charged polymers $P^{-}$, and vice versa for $e^{-}$. Consequently, the overall system is neutral, and excess ions play the role of salt, which could in principle screen electrostatic effects between polymers. Additional simulation details are described in the Supporting Text [20]. The snapshots of equilibrated MD simulations shown in Fig. 1b demonstrate that polymers form dense droplets surrounded by a dilute phase. Importantly, droplets are smaller for larger charge asymmetry $\Delta z=z_{P^{+}}+z_{P^{-}}$, indicating that charges can control droplet size and multiple droplets coexist stably.
@@
To reveal the physical mechanism of the observed droplet size regulation, we next describe the system using a continuous field theory. Here, the state of the incompressible, isothermal mixture is captured by the volume fraction fields $\phi_{P^{+}}(\boldsymbol{r})$, $\phi_{P^{-}}(\boldsymbol{r})$, $\phi_{\mathrm{e^{+}}}(\boldsymbol{r})$, and $\phi_{\mathrm{e^{-}}}(\boldsymbol{r})$ of the charged species. The remaining fraction, $\phi_{\mathrm{S}}=1-\sum_{i}\phi_{i}$ (here and below sums are over the four species $P^{+}$, $P^{-}$, $e^{+}$, and $e^{-}$), is filled by an inert solvent $S$. The system’s equilibrium is governed by the minimum of the total free energy $F=F_{\mathrm{FH}}+F_{\mathrm{int}}+F_{\mathrm{el}}$, where the terms respectively capture local, interfacial, and long-ranged electrostatic interactions akin to Ref. [11]. We approximate local interactions using the Flory–Huggins model [21, 22],
$F_{\mathrm{FH}}=\frac{k_{\mathrm{B}}T}{\nu}\!\int\bigg{[}\chi\phi_{P^{-}}\phi_{P^{+}}+\phi_{\mathrm{S}}\ln\phi_{\mathrm{S}}+\sum_{i}\frac{\phi_{i}}{l_{i}}\ln\phi_{i}\biggl{]}\mathrm{d}V\;,$ (1)
@@
where integrals are over the system of volume $V$, $k_{\mathrm{B}}T$ is the relevant thermal energy scale, and $\nu$ denotes the molecular volume, where solvent molecules, ions, and monomers of the polymers have the same volume for simplicity. In the integrand, the first term proportional to the Flory parameter $\chi<0$ captures the short-ranged attraction between $P^{+}$ and $P^{-}$, whereas the other terms capture translational entropies. To mimic the MD simulations, we chose the respective chain lengths $l_{i}$ as $l_{e^{+}}=l_{e^{-}}=1$ and $l_{P^{+}}=l_{P^{-}}=10$. The short-ranged interaction also implies an interfacial energy, which we describe by a simple gradient term [23]
@@
where $\varepsilon$ is the dielectric constant, assumed to be constant in space. Taken together, the equilibrium state thus depends on the local attraction ($\chi$), the interfacial penalty ($\kappa$), and the charge numbers ($z_{P^{+}}$ and $z_{P^{-}}$).
@@
Figure 2: Field theory predicts droplet-like periodic patterns. Periodic profiles of charge density $\rho_{i}=z_{i}\phi_{i}$ for all species $i$ and total charge density $\rho=\sum_{i}\rho_{i}$ as a function of position $x$ for weak (left) and strong (right) charge asymmetry $\Delta z=z_{P^{+}}+z_{P^{-}}$. Parameters are $z_{P^{-}}=-0.2$, $l_{P^{+}}=l_{P^{-}}=10$, $l_{e^{+}}=l_{e^{-}}=1$, $\bar{\phi}_{P^{+}}=\bar{\phi}_{P^{-}}=0.1$, $\kappa=10\,\ell_{\mathrm{B}}^{2}$, and $\chi=-5$, with Bjerrum length $\ell_{\mathrm{B}}=e^{2}/(\varepsilon k_{\mathrm{B}}T)$.
@@
As in the MD simulations, we impose constant average fractions $\bar{\phi}_{i}=V^{-1}\int\phi_{i}\mathrm{d}V$ with $\bar{\phi}_{P^{+}}=\bar{\phi}_{P^{-}}=0.1$, $\bar{\phi}_{e^{-}}=z_{P^{+}}\bar{\phi}_{P^{+}}$, and $\bar{\phi}_{e^{+}}=-z_{P^{-}}\bar{\phi}_{P^{-}}$. Based on the results shown in Fig. 1b, we expect that equilibrium states exhibit droplets of a well-defined size, which corresponds to periodic profiles in the field theory. To describe such states, we consider a one-dimensional, periodic system of variable length $L$, and minimize $F$ by varying $L$, $\psi(x)$, and $\phi_{i}(x)$. Here, we allow for coexisting phases with different periods and we impose mass conservation, incompressibility, and charge neutrality within each phase using Lagrange multipliers, akin to Ref. [25] and described in the Supporting Text [20]. The resulting equilibrium profiles shown in Fig. 2 indicate that the field theory indeed predicts periodic patterns, where regions of large polymer density correspond to the droplets and the surrounding region enriched in ions represent the dilute phase. Note that the ions do not neutralize the system everywhere (black line), which we confirmed in MD simulations (Fig. S1 in the Supporting Information [20]). These net charges are reminiscent of the double layer structure of complex coacervates [26] and indicate that electrostatics are crucial for droplet size control [27]. Indeed, we observe larger periods $L$ for smaller charge asymmetry (Fig. 2), consistent with the larger droplets in our MD simulations (Fig. 1b). There are also subtle differences between the two panels: For small charge asymmetry (Fig. 2a), the charge density $\rho$ exhibits a dip inside the droplet and reaches neutrality ($\rho=0$) in the dilute phase. Both effects are absent for larger charge asymmetry (Fig. 2b), suggesting that these two cases are qualitatively different.
@@
Figure 3: Field theory predicts transitions to patterned phases. (a) Phase diagram as a function of the charge numbers $z_{P^{+}}$ and $z_{P^{-}}$ of the polymers revealing parameter regions with the coexistence of two homogeneous phases (H+H) and patterned phases (P) for strong attraction ($\chi=-5$). The colored markers correspond to the two panels in Fig. 2. (b) Phase diagram for weak attraction ($\chi=-4.5$) including regions with a single homogeneous phase (H). The remaining columns show respective periods $L$ and amplitudes $A_{P^{+}}=\max(\phi_{P^{+}})-\min(\phi_{P^{+}})$. (c) Schematic explanation of the origin of the patterned phase. The initial phase-separated charge-neutral state in (i) is destabilized due to net charges if charge asymmetry is enabled in (ii). An equilibrium state can either be reached by restoring charge neutrality in (iii) or by breaking the phases into smaller droplets in (iv). (d) Pattern length scale $L$ as a function of interaction parameter $\chi$ (left, interfacial parameter $\kappa=10$) and $\kappa$ (right, various $\chi$) for various $z_{P^{+}}$ at $z_{P^{-}}=-0.2$. As $|\chi|$ decreases, the patterned phase transitions to macrophase separation (stars) or a homogeneous state (open disk). Dashed lines indicate the most unstable length scale predicted by the linear stability analysis in the Supporting Text [20]. (a–d) Additional parameters are given in Fig. 2.
@@
We next analyze in detail how the periodic patterns depend on the charge numbers $z_{P^{+}}$ and $z_{P^{-}}$. Fig. 3a shows that periodic patterns only appear in a restricted parameter region (yellow region P). In particular, this patterned phase does not exist for symmetric mixtures (diagonal line, $\Delta z=0$). Instead, the system exhibits macrophase separation (blue H+H region) if $\chi$ is sufficiently negative, and this region becomes larger for weaker interactions (Fig. 3b and Fig. S21 [20]). Moreover, the system stays homogeneous for very asymmetric mixtures (teal H region). Collectively, these observations suggest periodic patterns emerge for sufficiently strong attraction (large $-\chi$) and intermediate charge asymmetry $|\Delta z|$ between polymers.
@@
We can understand the emergence of patterned phases starting from the charge-balanced state ($\Delta z=0$, Fig. 3c, subpanel i). In this case, the polymers form a dense macrophase while small species ($e^{+}$, $e^{-}$, and $S$) accumulate in the corresponding dilute phase. Here, we have $F=F_{\mathrm{FH}}$ since $F_{\mathrm{el}}=0$ and the energy $F_{\mathrm{int}}$ of the single interface is negligible in a thermodynamically large system. Switching on charge asymmetry ($\Delta z\neq 0$, subpanel ii) destabilizes this state because the macrophases acquire a net charge, so the electrostatic energy $F_{\mathrm{el}}$ diverges. To minimize the total free energy $F$, the state can change in two fundamentally different ways: it can either restore charge neutrality (subpanel iii) or form a patterned phase (subpanel iv). In the first alternative, ions move into the dense phase (maintaining macrophase separation, subpanel iii), which restores charge neutrality and maintains $F_{\mathrm{el}}=F_{\mathrm{int}}=0$. However, the ions decrease the concentration of polymers, thus reducing short-ranged contacts between polymers, and increasing $F_{\mathrm{FH}}$. In the second alternative, the dense phase is broken up to form a pattern of finite length scale $L$ (subpanel iv). In the simplest case, the polymers maintain a high concentration, implying that $F_{\mathrm{FH}}$ hardly changes. However, the electrostatic energy $F_{\mathrm{el}}$ is now finite, and the breakup creates many interfaces, so both $F_{\mathrm{el}}$ and $F_{\mathrm{int}}$ are positive. Taken together, both alternatives increase $F$ compared to the neutral initial case, but breakup into the patterned phase is favored for strong attraction (large $-\chi$).
@@
The preceding argument also makes qualitative predictions for the pattern period $L$ as a proxy for droplet size. Assuming that the density of polymers stays constant in the patterned phase, $F_{\mathrm{FH}}$ is independent of $L$, so $L$ is instead governed by the minimum of $F_{\mathrm{int}}+F_{\mathrm{el}}$. Generally, $F_{\mathrm{int}}$ decreases with increasing $L$ (since there are fewer interfaces per unit length), whereas $F_{\mathrm{el}}$ increases with $L$ (since opposite charges are separated further for larger $L$, i.e., exchanging the two central slabs in Fig. 3c(iv) costs electrostatic energy). These arguments predict that $L$ increases with larger $\kappa$ (increasing $F_{\mathrm{int}}$), whereas $L$ decreases with larger $|\Delta z|$ (increasing $F_{\mathrm{el}}$). We present a more quantitative argument in the Supporting Text [20], which shows that at the free energy minimum the two energies have opposing dependencies on $L$, $\mathrm{d}F_{\mathrm{int}}/\mathrm{d}L=-\mathrm{d}F_{\mathrm{el}}/\mathrm{d}L$, leading to a trade-off where $F_{\mathrm{int}}=F_{\mathrm{el}}$ at the minimum. We thus generally expect that $L$ increases with larger $\kappa$ and smaller $|\Delta z|$.
@@
We test our predictions by analyzing the patterned phases obtained from the field theory. Fig. 3b shows that $L$ indeed decreases for larger $|\Delta z|$ (away from the symmetry line), although the patterns completely vanish for large $|\Delta z|$ where the homogeneous state exhibits the lowest free energy. The right-most panel in Fig. 3b shows that the amplitude of $\phi_{P^{+}}(x)$ decreases with increasing $|\Delta z|$, implying that polymers become less concentrated. This is not captured by the simple argument above, suggesting that the system compromises the segregation of polymers (implying larger $F_{\mathrm{FH}}$) for a lower electrostatic energy $F_{\mathrm{el}}$. Indeed we find that $F_{\mathrm{el}}$ exhibits a non-monotonous dependency on $|\Delta z|$ (Supporting Figure S11 [20]), and the patterns cease when $F_{\mathrm{el}}$ vanishes. Concomitantly, the amplitude approaches zero (Fig. 3b), suggesting that the transition to the homogeneous phase is continuous.
@@
To also test the influence of the interaction strength $\chi$ and the interfacial parameter $\kappa$ on the pattern size $L$, we performed additional minimizations. Fig. 3d confirms a weak influence of $\chi$, whereas $L$ increases significantly with $\kappa$. Note that these length scales cannot be predicted from a simple stability analysis of the homogeneous state (dashed line in Fig. 3d; details in the Supporting Text [20]), except at the transition of the patterned phase to the homogeneous state (open circles). This observation, and the fact the scaling is consistent with mean-field theory (Supporting Figure S5 [20]), provides additional evidence for a continuous phase transition. Taken together, the numerical solutions of the field theory confirm that $L$ (and thus the droplet size) increases with larger $\kappa$ and smaller $|\Delta z|$, while $\chi$ has a minor influence beyond driving droplet formation in the first place.
@@
Figure 4: MD simulations corroborate field theory. (a) Simulations at various charge numbers ($z_{P^{+}}$, $z_{P^{-}}$) demonstrate that high charge asymmetries result in smaller average droplet sizes (normalized by total number of polymer chains in the system). (b) Normalized droplet size as a function of attraction strength $\epsilon_{\mathrm{LJ}}$. Decreasing $\epsilon_{\mathrm{LJ}}$ mildly increases or decreases droplet size depending on charge asymmetry.